### v1

@@ 4 Preparatory results
$\begin{split}|T_{2}|&\leq{}\bigg{\{}\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}+\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!h_{F}^{-1}\|{\gamma_{K(F),F}^{\mathrm{c}}({\boldsymbol{v}})}\|_{F}^{2}\bigg{\}}^{\frac{1}{2}}\bigg{\{}\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}\cup\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!h_{F}\|{{\boldsymbol{\phi}}}\|_{F}^{2}\bigg{\}}^{\frac{1}{2}}\\ &\lesssim h_{{\mathbf{a}}}\bigg{\{}\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}+\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!h_{F}^{-1}\|{\gamma_{K(F),F}^{\mathrm{c}}({\boldsymbol{v}})}\|_{F}^{2}\bigg{\}}^{\frac{1}{2}}\|{{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}}\|_{D_{{\mathbf{a}}}}.\end{split}$
@@
For mixed boundary conditions, Lemma 4.3 is adapted as follows: There exist a constant $C(\kappa_{\mathcal{T}_{h}})$ so that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$ and all ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$, there exists ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{\boldsymbol{H}}_{\partial D_{{\mathbf{a}}}\backslash\Gamma_{\rm N}}(\text{\bf curl};D_{{\mathbf{a}}})$ satisfying ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{\boldsymbol{H}}_{\partial D_{{\mathbf{a}}}\cap\Gamma_{\rm N}}(\text{\rm div}=0;D_{{\mathbf{a}}})$ such that
@@
To prove this inequality, one adapts the above proof as follows: First, [2, Lemma 3.5] can be replaced by [28, Theorem 3] on regular decompositions with mixed boundary conditions, but the stability constant is not given explicitly in terms of the geometry of $D_{{\mathbf{a}}}$. To have a stability constant only depending on the mesh shape-regularity, one needs to conjecture that the result of [27] can be extended to mixed boundary conditions. We notice that the extension holds for pure Neumann conditions, as shown in [27, Theorem 32].
@@
One can adapt the proof of Lemma 4.3 to establish the following local broken-divergence, curl-preserving Poincaré inequality: There exists a constant $C(\kappa_{\mathcal{T}_{h}})$ so that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$ and all ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$, there exist ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{{{\boldsymbol{H}}_{0}(\text{\rm div};D_{{\mathbf{a}}})}}$ satisfying ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\bf curl}={\boldsymbol{0}};D_{{\mathbf{a}}})}}$ such that,
$\|{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\|_{D_{{\mathbf{a}}}}\!\!\leq\!{}C(\kappa_{\mathcal{T}_{h}})h_{{\mathbf{a}}}\bigg{\{}\!\|\nabla_{h}{\cdot}({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}})\|_{D_{{\mathbf{a}}}}^{2}\!+\!\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{d}}_{F}}\|_{F}^{2}+\!\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!\!h_{F}^{-1}\|{\gamma_{K(F),F}^{\mathrm{d}}({\boldsymbol{v}})}\|_{F}^{2}\!\bigg{\}}^{\frac{1}{2}}.$
@@
Under Assumption 3.5, there exists a constant $C(\kappa_{\mathcal{T}_{h}})$ such that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$ and all ${\boldsymbol{v}}\in{\boldsymbol{X}^{\mathrm{c}}}(D_{{\mathbf{a}}})$ (see (3.2b)),
$|{\boldsymbol{v}}|_{{\boldsymbol{H}}^{\frac{1}{2}+\delta}(K)}\leq C(\kappa_{\mathcal{T}_{h}})h_{{\mathbf{a}}}^{\frac{1}{2}-\delta}\|{\nabla{\times}{\boldsymbol{v}}}\|_{D_{{\mathbf{a}}}}\quad\forall K\in{\cal{T}}_{{\mathbf{a}}},$ (4.5)
where the real number $\delta:=\delta(\kappa_{\mathcal{T}_{h}})\in(0,\frac{1}{2}]$ comes from Assumption 3.5.
@@
(1) Let ${\mathbf{a}}\in\mathcal{V}_{h}$ and ${\boldsymbol{v}}\in{\boldsymbol{X}^{\mathrm{c}}}(D_{{\mathbf{a}}})$. Invoking the same arguments as in Step (2) of the proof of Lemma 4.3, we infer that there exists ${\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}})$ such that $\nabla{\times}{\boldsymbol{\phi}}=\nabla{\times}{\boldsymbol{v}}$ and $\nabla{\cdot}{\boldsymbol{\phi}}=0$ in $D_{{\mathbf{a}}}$, and
@@
Since $\nabla{\times}({\boldsymbol{v}}-{\boldsymbol{\phi}})={\boldsymbol{0}}$, there exists a function $p\in X^{\mathrm{g}}(D_{{\mathbf{a}}})$ such that $\nabla p={\boldsymbol{v}}-{\boldsymbol{\phi}}$. Altogether, this leads to the regular decomposition
${\boldsymbol{v}}=\nabla p+{\boldsymbol{\phi}},\quad p\in X^{\mathrm{g}}(D_{{\mathbf{a}}}),\quad{\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}}).$
(2) Since ${\boldsymbol{v}}\in{{{\boldsymbol{H}}_{0}(\text{\rm div}=0;D)}}$ by assumption, we infer that $(\nabla p,\!\nabla\rho)_{D_{{\mathbf{a}}}}\!\!=\!({\boldsymbol{\phi}},\!\nabla\rho)_{D_{{\mathbf{a}}}}$ for all $\rho\in X^{\mathrm{d}}(D_{{\mathbf{a}}})$. Therefore, we can apply Assumption 3.5 to ${\boldsymbol{g}}:={\boldsymbol{\phi}}$, and observing that $\zeta_{{\boldsymbol{g}}}=p$, we infer that
@@
(3) For all $K\in{\cal{T}}_{{\mathbf{a}}}$, invoking the geometric mapping to the reference simplex $^{\widehat{K}}$ and the embedding $|\widehat{\boldsymbol{\phi}}|_{{\boldsymbol{H}}^{\frac{1}{2}}({\widehat{K}})}\leq{\widehat{C}}\|\nabla\widehat{\boldsymbol{\phi}}\|_{{\widehat{K}}}$, we infer that
@@ 5 Proof of the main results from Section 3
This section contains the proofs of Lemmas 3.3, 3.4, and 3.6.
@@
Invoking the Cauchy–Schwarz inequality, the stability bound (4.1b) and the fact that $\psi_{{\mathbf{a}}}$ is continuous and bounded by one, we infer that
@@
where we used (4.1b) in the last bound.
@@
$\mbox{II}=(\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}),\nabla(\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi)))_{D_{{\mathbf{a}}}}+(\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}),\nabla{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi))_{D_{{\mathbf{a}}}}=:\mbox{II}_{1}+\mbox{II}_{2},$
@@
$\mbox{II}_{1}=\sum_{K\in{\cal{T}}_{{\mathbf{a}}}}(\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi),{\boldsymbol{n}}_{K}{\cdot}(\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})))_{\partial K}.$
All the mesh faces in $\bigcup_{K\in{\cal{T}}_{{\mathbf{a}}}}{\mathcal{F}}_{K}$ are either in $\mathcal{F}_{{\mathbf{a}}}$ or in $\mathcal{F}_{{\mathbf{a}}}^{\partial}$. In the first case, we observe that $\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi)$ is single-valued across every interface $F\in\mathcal{F}_{{\mathbf{a}}}\cap{\mathcal{F}}_{h}^{\mathrm{int}}$ and that we have set conventionally ${\bullet}]\!]^{\mathrm{d}}_{F}:=\bullet|_{F}{\cdot}{\boldsymbol{n}}_{D}$ on every boundary face $F\in\mathcal{F}_{{\mathbf{a}}}\cap{\mathcal{F}}_{h}^{\mathrm{bd}}$. In the second case, $\psi_{{\mathbf{a}}}$ vanishes identically on the face in question, and so does the normal component of $\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})$. Altogether, this gives
$\mbox{II}_{1}={}$ $\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}(\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi),[\![{\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})}]\!]^{\mathrm{d}}_{F})_{F}.$
@@
$|\mbox{II}_{12}|\lesssim{}$ $p^{-\frac{3}{2}}\bigg{\{}\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!\Big{(}\frac{p^{2}}{h_{F}}\Big{)}\|{[\![{{\boldsymbol{E}}_{h}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}\bigg{\}}^{\frac{1}{2}}\|{\nabla_{h}{\times}{\boldsymbol{\delta}}_{{\mathbf{a}}}}\|_{D_{{\mathbf{a}}}}.$
@@
(3) Putting the bounds on $\mbox{II}_{1}$ and $\mbox{II}_{2}$ together and since $p\geq 1$, we conclude the proof.
@@
We apply the local broken-curl, divergence-preserving Poincaré inequality to ${\boldsymbol{v}}:=\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}\in{\boldsymbol{P}}^{\mathrm{b}}_{p+1}({\cal{T}}_{{\mathbf{a}}})\subset{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$ and select ${\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}:={\boldsymbol{E}}_{{\mathbf{a}}}$, which indeed satisfies ${\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ and ${\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}-{\boldsymbol{v}}=\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}-{\boldsymbol{U}}_{\mathbf{a}}-\nabla\theta_{\mathbf{a}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ since ${\boldsymbol{U}}_{\mathbf{a}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ by construction and $\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}-\nabla\theta_{\mathbf{a}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ by definition of $\theta_{\mathbf{a}}$. Observing that ${\boldsymbol{\delta}}_{{\mathbf{a}}}={\boldsymbol{E}}_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}={\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}-{\boldsymbol{v}}$, we infer that
$\|{\boldsymbol{\delta}}_{{\mathbf{a}}}\|_{D_{{\mathbf{a}}}}$ $\lesssim h_{\mathbf{a}}\bigg{\{}\|\nabla_{h}{\times}{\boldsymbol{\delta}}_{\mathbf{a}}\|_{D_{{\mathbf{a}}}}+\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!h_{F}^{-1}\|{[\![{\psi_{\mathbf{a}}{\boldsymbol{E}}_{h}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}+\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!h_{F}^{-1}\|{\gamma_{K(F),F}^{\mathrm{c}}(\psi_{\mathbf{a}}{\boldsymbol{E}}_{h})}\|_{F}^{2}\bigg{\}}^{\frac{1}{2}}$
@@
This problem is well-posed since $D_{{\mathbf{a}}}$ has trivial topology and ${\boldsymbol{\delta}}_{{\mathbf{a}}}$ is divergence-free. Indeed, $\nabla{\cdot}{\boldsymbol{\delta}}_{{\mathbf{a}}}=\nabla{\cdot}({\boldsymbol{E}}_{{\mathbf{a}}}-(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}))=\nabla{\cdot}{\boldsymbol{U}}_{{\mathbf{a}}}+\nabla{\cdot}(\nabla\theta_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})=0$ since ${\boldsymbol{U}}_{{\mathbf{a}}}$ is divergence-free by construction and the definition (3.5) of $\theta_{{\mathbf{a}}}$ implies that $\nabla\theta_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}$ is divergence-free as well. Recalling the regularity pickup from [16, Section 5.2], and [23, Lemma 44.2], we infer that there is $\delta_{\mathbf{a}}\in(0,\frac{1}{2}]$ (here, $\delta_{\mathbf{a}}$ can depend on ${\mathbf{a}}\in\mathcal{V}_{h}$) so that $\nabla_{\!0}{\times}{\boldsymbol{Z}}_{{\mathbf{a}}}\in{\boldsymbol{H}}^{\frac{1}{2}+\delta}(D_{{\mathbf{a}}})$. This implies in particular that $\nabla_{\!0}{\times}{\boldsymbol{Z}}_{{\mathbf{a}}}$ has meaningful (tangential) traces in ${\boldsymbol{L}}^{2}(\partial K)$ for all $K\in{\cal{T}}_{{\mathbf{a}}}$. Taking the ${\boldsymbol{L}}^{2}(D_{{\mathbf{a}}})$-inner product with ${\boldsymbol{\delta}}_{{\mathbf{a}}}$ in (5.3a) and using the definition of ${\boldsymbol{\delta}}_{{\mathbf{a}}}$, we obtain
@@
Next, invoking the relation (3.4) with ${\boldsymbol{W}}_{{\mathbf{a}}}:={\boldsymbol{Z}}_{{\mathbf{a}}}\in{\boldsymbol{X}^{\mathrm{c}}_{0}}(D_{{\mathbf{a}}})$ and the definition (2.9) of lifting operator, we infer that

### v2

@@ 4 Preparatory results
$\begin{split}|T_{2}|&\leq{}\bigg\{\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}\bigg\}^{\frac{1}{2}}\bigg\{\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!h_{F}\|{{\boldsymbol{\phi}}}\|_{F}^{2}\bigg\}^{\frac{1}{2}}\\ &\lesssim h_{{\mathbf{a}}}\bigg\{\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}\bigg\}^{\frac{1}{2}}\|{{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}}\|_{D_{{\mathbf{a}}}}.\end{split}$
@@
For mixed boundary conditions, the statement of Lemma 4.19 is adapted as follows: There exists a constant $C(\kappa_{\mathcal{T}_{h}})$ so that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$, all ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$, and all ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{\boldsymbol{H}}_{\partial D_{{\mathbf{a}}}\backslash\Gamma_{\rm N}}(\text{\bf curl};D_{{\mathbf{a}}})$ with ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{\boldsymbol{H}}_{\partial D_{{\mathbf{a}}}\cap\Gamma_{\rm N}}(\text{\rm div}=0;D_{{\mathbf{a}}})$, the following holds:
@@
The proof of this inequality is, however, not straightforward. First, [2, Lemma 3.5] can be replaced by [29, Theorem 3] on regular decompositions with mixed boundary conditions, but the stability constant is not given explicitly in terms of the geometry of $D_{{\mathbf{a}}}$. To have a stability constant only depending on the mesh shape-regularity, one needs to conjecture that the result of [28] can be extended to mixed boundary conditions. We notice that the extension holds for pure Neumann conditions [28, Theorem 32].
@@
One can adapt the proof of Lemma 4.19 to establish the following local broken-divergence, curl-preserving Poincaré inequality: There exists a constant $C(\kappa_{\mathcal{T}_{h}})$ so that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$, all ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$, and all ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{{{\boldsymbol{H}}_{0}(\text{\rm div};D_{{\mathbf{a}}})}}$ with ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\bf curl}={\boldsymbol{0}};D_{{\mathbf{a}}})}}$, the following holds:
$\|{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\|_{D_{{\mathbf{a}}}}\leq{}C(\kappa_{\mathcal{T}_{h}})h_{{\mathbf{a}}}\bigg\{\|\nabla_{h}{\cdot}({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}})\|_{D_{{\mathbf{a}}}}^{2}+\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{d}}_{F}}\|_{F}^{2}\bigg\}^{\frac{1}{2}}.$
@@
Under Assumption 3.9 with $\delta:=\delta(\kappa_{\mathcal{T}_{h}})\in(0,\frac{1}{2}]$, there exists a constant $C(\kappa_{\mathcal{T}_{h}})$ (depending on $\delta$) such that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$ and all ${\boldsymbol{v}}\in{\boldsymbol{X}^{\mathrm{c}}}(D_{{\mathbf{a}}})$ (see (3.2b)),
$|{\boldsymbol{v}}|_{{\boldsymbol{H}}^{\frac{1}{2}+\delta}(K)}\leq C(\kappa_{\mathcal{T}_{h}})h_{{\mathbf{a}}}^{\frac{1}{2}-\delta}\|{\nabla{\times}{\boldsymbol{v}}}\|_{D_{{\mathbf{a}}}}\quad\forall K\in{\cal{T}}_{{\mathbf{a}}}.$ (4.4)
@@
(1) Let ${\mathbf{a}}\in\mathcal{V}_{h}$ and ${\boldsymbol{v}}\in{\boldsymbol{X}^{\mathrm{c}}}(D_{{\mathbf{a}}})$. Invoking the same arguments as in Step (1) of the proof of Lemma 4.19, we infer that there exists ${\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}})$ such that $\nabla{\times}{\boldsymbol{\phi}}=\nabla{\times}{\boldsymbol{v}}$ in $D_{{\mathbf{a}}}$, and
@@
Since $\nabla{\times}({\boldsymbol{v}}-{\boldsymbol{\phi}})={\boldsymbol{0}}$, there exists a function $p\in X^{\mathrm{g}}(D_{{\mathbf{a}}})$ such that $\nabla p={{\boldsymbol{\phi}}-{\boldsymbol{v}}}$. Altogether, this leads to the regular decomposition
${\boldsymbol{v}}={-}\nabla p+{\boldsymbol{\phi}},\quad p\in X^{\mathrm{g}}(D_{{\mathbf{a}}}),\quad{\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}}).$
(2) Since ${\boldsymbol{v}}\in{{{\boldsymbol{H}}_{0}(\text{\rm div}=0;D)}}$ by assumption, we infer that $(\nabla p,\!\nabla\rho)_{D_{{\mathbf{a}}}}\!\!=\!({\boldsymbol{\phi}},\!\nabla\rho)_{D_{{\mathbf{a}}}}$ for all $\rho\in{X^{\mathrm{g}}}(D_{{\mathbf{a}}})$. Therefore, we can apply Assumption 3.9 to ${\boldsymbol{g}}:={\boldsymbol{\phi}}$, and observing that $\zeta_{{\boldsymbol{g}}}=p$, we infer that
@@
(3) For all $K\in{\cal{T}}_{{\mathbf{a}}}$, invoking the geometric mapping to the reference simplex $^{\widehat{K}}$ and the embedding $|\widehat{\boldsymbol{\phi}}|_{{\boldsymbol{H}}^{\frac{1}{2}{+\delta}}({\widehat{K}})}\leq{\widehat{C}}\|\nabla\widehat{\boldsymbol{\phi}}\|_{{\widehat{K}}}$, we infer that
@@ 5 Proof of the main results from Section 3
This section contains the proofs of Lemmas 3.7, 3.8, and 3.10.
@@
Invoking the Cauchy–Schwarz inequality and the fact that $\psi_{{\mathbf{a}}}$ is continuous and bounded by one, we infer that
@@
where we used the stability bound (4.1b) in the last estimate.
@@
$\mbox{II}=-(\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}),\nabla(\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi)))_{D_{{\mathbf{a}}}}-(\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}),\nabla{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi))_{D_{{\mathbf{a}}}}=:\mbox{II}_{1}+\mbox{II}_{2},$
@@
$\mbox{II}_{1}=-\sum_{K\in{\cal{T}}_{{\mathbf{a}}}}(\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi),{\boldsymbol{n}}_{K}{\cdot}(\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})))_{\partial K}.$
All the mesh faces in $\bigcup_{K\in{\cal{T}}_{{\mathbf{a}}}}{\mathcal{F}}_{K}$ are either in $\mathcal{F}_{{\mathbf{a}}}$ or in ${\mathcal{F}}_{\setminus{\boldsymbol{a}}}$. In the first case, we observe that $\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi)$ is single-valued across every interface $F\in\mathcal{F}_{{\mathbf{a}}}\cap{\mathcal{F}}_{h}^{\mathrm{int}}$ and that we have set conventionally ${\bullet}]\!]^{\mathrm{d}}_{F}:=\bullet|_{F}{\cdot}{\boldsymbol{n}}_{D}$ on every boundary face $F\in\mathcal{F}_{{\mathbf{a}}}\cap{\mathcal{F}}_{h}^{\mathrm{bd}}$. In the second case where $F\in{\mathcal{F}}_{\setminus{\boldsymbol{a}}}$, we have ${\boldsymbol{n}}_{F}{\cdot}\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})={\boldsymbol{n}}_{F}{\cdot}(\psi_{\mathbf{a}}{\nabla_{h}{\times}{\boldsymbol{E}}_{h}})+{{\boldsymbol{E}}_{h}}{\cdot}({\boldsymbol{n}}_{F}{\times}\nabla\psi_{\mathbf{a}})=0$ since $\psi_{{\mathbf{a}}}=0$ and $({\boldsymbol{n}}_{F}{\times}\nabla\psi_{{\mathbf{a}}})_{F}={\boldsymbol{0}}$ on $F$. Altogether, this gives
$\mbox{II}_{1}={}$ $-\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}(\xi-{{\mathcal{I}}}^{{\rm g}}_{{\mathbf{a}}}(\xi),[\![{\nabla_{h}{\times}(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})}]\!]^{\mathrm{d}}_{F})_{F}.$
@@
$|\mbox{II}_{12}|\lesssim{}$ $\bigg\{\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!p^{-3}\Big(\frac{p^{2}}{h_{F}}\Big)\|{[\![{{\boldsymbol{E}}_{h}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}\bigg\}^{\frac{1}{2}}\|{\nabla_{h}{\times}{\boldsymbol{\delta}}_{{\mathbf{a}}}}\|_{D_{{\mathbf{a}}}}.$
@@
(3) Putting the bounds on $\mbox{II}_{1}$ and $\mbox{II}_{2}$ together and since $p\geq 1$, we conclude the proof. (Notice that we just drop the factor $p^{-3}$ in the bound on $\mbox{II}_{1}$.)
@@
We apply the local broken-curl, divergence-preserving Poincaré inequality to ${\boldsymbol{v}}:=\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}\in{\boldsymbol{P}}^{\mathrm{b}}_{p+1}({\cal{T}}_{{\mathbf{a}}})\subset{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$ and select ${\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}:={\boldsymbol{E}}_{{\mathbf{a}}}$, which indeed satisfies ${\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ and ${\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}-{\boldsymbol{v}}={\boldsymbol{U}}_{\mathbf{a}}+\nabla\theta_{\mathbf{a}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ since ${\boldsymbol{U}}_{\mathbf{a}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ by construction and $\nabla\theta_{\mathbf{a}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ by definition of $\theta_{\mathbf{a}}$. Observing that ${\boldsymbol{\delta}}_{{\mathbf{a}}}={\boldsymbol{E}}_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}={\boldsymbol{v}}^{\mathrm{c}}_{\mathbf{a}}-{\boldsymbol{v}}$, we infer that
$\|{\boldsymbol{\delta}}_{{\mathbf{a}}}\|_{D_{{\mathbf{a}}}}$ $\lesssim h_{\mathbf{a}}\bigg\{\|\nabla_{h}{\times}{\boldsymbol{\delta}}_{\mathbf{a}}\|_{D_{{\mathbf{a}}}}+\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!h_{F}^{-1}\|{[\![{\psi_{\mathbf{a}}{\boldsymbol{E}}_{h}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}\bigg\}^{\frac{1}{2}}$
@@
This problem is well-posed since $D_{{\mathbf{a}}}$ has trivial topology and ${\boldsymbol{\delta}}_{{\mathbf{a}}}$ is divergence-free. Indeed, $\nabla{\cdot}{\boldsymbol{\delta}}_{{\mathbf{a}}}=\nabla{\cdot}({\boldsymbol{E}}_{{\mathbf{a}}}-(\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}))=\nabla{\cdot}{\boldsymbol{U}}_{{\mathbf{a}}}+\nabla{\cdot}(\nabla\theta_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h})=0$ since ${\boldsymbol{U}}_{{\mathbf{a}}}$ is divergence-free by construction and the definition (3.5) of $\theta_{{\mathbf{a}}}$ implies that $\nabla\theta_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}$ is divergence-free as well. Recalling the regularity pickup from [16, Section 5.2], and [23, Lemma 44.2], we infer that there is ${\delta_{\mathbf{a}}^{\prime}}\in(0,\frac{1}{2}]$ (here, $\delta_{\mathbf{a}}^{\prime}$} can depend on ${\mathbf{a}}\in\mathcal{V}_{h}$) so that $\nabla_{\!0}{\times}{\boldsymbol{Z}}_{{\mathbf{a}}}\in{\boldsymbol{H}}^{\frac{1}{2}+{\delta_{\mathbf{a}}^{\prime}}}(D_{{\mathbf{a}}})$. This implies in particular that $\nabla_{\!0}{\times}{\boldsymbol{Z}}_{{\mathbf{a}}}$ has meaningful (tangential) traces in ${\boldsymbol{L}}^{2}(\partial K)$ for all $K\in{\cal{T}}_{{\mathbf{a}}}$. Taking the ${\boldsymbol{L}}^{2}(D_{{\mathbf{a}}})$-inner product with ${\boldsymbol{\delta}}_{{\mathbf{a}}}={\boldsymbol{U}}_{{\mathbf{a}}}+\nabla\theta_{{\mathbf{a}}}-\psi_{{\mathbf{a}}}{\boldsymbol{E}}_{h}$ in (5.3a), we obtain
@@
Next, invoking the relation (3.4) with ${\boldsymbol{W}}_{{\mathbf{a}}}:={\boldsymbol{Z}}_{{\mathbf{a}}}\in{\boldsymbol{X}^{\mathrm{c}}_{0}}(D_{{\mathbf{a}}})$ and the definition (2.9) of the lifting operator, we infer that